If P(E)=0.59, P(F)=0.61, and P(E and F)=0.83 find P(E and F)

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

P(E and F)=P(E)+P(F)-P(E OR F) rearranging the addition property of probabilities and making the assumption that p(E and F) is truly p(E OR F), otherwise the problem doesn't make sense (and can't be more than each; or can be).
p(E and F)=0.59+0.61-0.83=0.37

Answer by ikleyn(52754)

.
Your post is wrong.